2.5kg of water is heated in an open container for 5 minutes using a heater of power 2KW. Which of the following statements about the temperature of the water is true?(the SHC of water= 4200Jkg-1k-1)


A. The change in temperature of the water is 58.30C


B. The change in temperature of the water is 57.30C


C. The change in temperature of the water is greater 57.30C


D.If the initial temperature of the water was 23.70C then its final temperature would be 8100C


1
Expert's answer
2021-08-17T10:30:07-0400

Solution.

"m=2.5kg;"

"t=5min=300s;"

"P=2000W;"

"c=4200Jkg^{-1}K^{-1};"

"A=Pt;"

"Q=cm\Delta T;"

"Q=A;"

"Pt=cm\Delta T\implies \Delta T=\dfrac{Pt}{cm};"

"\Delta T=\dfrac{2000\sdot300}{4200\sdot2.5}=57.3K=57.3^oC;"

Answer:B. The change in temperature of the water is 57.3oC.